OpenStudy (anonymous):

can someone help me w/ this problem f(t)= 5000/1+50e^-0.4t a. approximately how many people will have heard the news after 8 days? b. how long will it take for half of the townspeople to have heard about the meeting? c. find an equation to express the rate at which notice of the meeting is spreading

OpenStudy (experimentx):

f(t)= 5000/(1+50e^-0.4t)

OpenStudy (callisto):

is the unit of t day?

OpenStudy (anonymous):

= 134

OpenStudy (anonymous):

@uh_oh_boy23 I believe your post's sort of chop off head and tail!

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

yes expx

OpenStudy (anonymous):

heres the rest Notice of a town meeting is spread by word of mouth to the town's 5000 residents. After t days, the number of people who know about the meeting is expressed by the function

OpenStudy (anonymous):

the beginning

OpenStudy (anonymous):

b) Half of town ...? There's no data about the number of the whole town?

OpenStudy (experimentx):

for your first proble, put 8 inplace of t and calculate it's vlaue

OpenStudy (experimentx):

i think the whole town people is 5000

OpenStudy (callisto):

f(8)= 5000/(1+50e^-0.4(8)) = 5000 / (1+50e^-3.2) Can you do the calculation yourself?

OpenStudy (callisto):

@experimentX it's 5000 as '' the town's 5000 residents''

OpenStudy (callisto):

(b) f(t) = 5000/2 = 5000/(1+50e^-0.4t) 2500 = 5000/(1+50e^-0.4t) 2 = (1+50e^-0.4t) Can you solve it from here?

OpenStudy (experimentx):

yeah .. my language mistake

OpenStudy (experimentx):

i think for the last part ... you just need to differentiate the function with respect to t

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so for like A plug in 8 for t?

OpenStudy (anonymous):

c) f'(t) = 10.000 e^ (- .04t) / [ 1 + 50 e (- .04t) ]²

OpenStudy (anonymous):

Why don't you give a try on a) and b)